4. A fluid moves in a steady flow manner between two sections in the same



flowline. At section 1: A = 0.10 m2, V, = 6 m/s, v1 = 0.33 m /kg. At



section 2: A2 = 0.2 m2, P2 = 0.27 kg/mº. Calculate for the velocity of flow



of section 2.



A. 37.63 m/s



C. 40.62 m/s



B. 33.67 m/s*



D. 46.02 m/s

1
Expert's answer
2021-12-22T14:13:02-0500

"\frac{A_1v_1}{V_1}=\rho_2A_2v_2,\implies"

"v_2=\frac{A_1v_1}{V_1\rho_2A_2}=33.67~\frac ms."